Tharaldson Corporation makes a product with the following standard costs: 


Standard
Standard

Quantity orStandard Price orCost Per

HoursRateUnit
Direct materials6.0 ounces$3.00 per ounce$18.00
Direct labor0.3 hours$11.00 per hour$3.30
Variable overhead0.3 hours$9.00 per hour$2.70


The company reported the following results concerning this product in June. 

Originally budgeted output3,600 units
Actual output3,200 units
Raw materials used in production21,000 ounces
Purchases of raw materials22,100 ounces
Actual direct labor-hours500 hours
Actual cost of raw materials purchases$42,300
Actual direct labor cost$13,600
Actual variable overhead cost$3,800

The company applies variable overhead on the basis of direct labor-hours. The direct materials purchases variance is computed when the materials are purchased. 

The labor efficiency variance for June is:
Answer #1

>> Labor Efficiency variance = ( Standard Hours - Actual Hours ) * Standard Rate.

>> Standard Hours = ( 3,200 * 0.3 ) = 960 Hours.

>> Labor Efficiency Variance = ( 960 - 500 ) * $ 11

>> Labor Efficiency Variance = $ 5.060 Favorable.